OpenStudy (anonymous):

find the unknown coordinate so the line through the points has the given slope (-3,2), (0,y); slope -2

OpenStudy (anonymous):

\[\frac{y-2}{0-(-3)}=-2\] or \[\frac{y-2}{3}=-2\] solve for \(y\)